### v1

# TIF-UNIMI-2024-12 Asymptotic Freedom in Parton Language: the Birth

Abstract
I review the contributions of Giorgio Parisi to perturbative QCD. Concentrated in a decade, they mark the transition of the theory of strong interactions from a set of loosely connected ideas based on models, to a quantum field theory that is now an integral part of the standard model of fundamental interactions. Parisi’s contributions have established at a very early stage ideas, methods and tools that are now standard, and in several cases anticipated results that only became prominent in the XXIst century.
Contribution to the volume
From Quantum Fields to Spin
Glasses:
A journey through the contributions of Giorgio Parisi to theoretical Physics
1 Parisi and QCD
The contributions of Giorgio Parisi to the theory of strong interactions in the perturbative domain are concentrated in one decade: the first paper [1] was published in 1970, and the last [2] in 1980. This is the decade of the “triumph of quantum field theory”, in Sidney Coleman’s words [3], filled with “wonderful things brought back from far places to make the spectator gasp with awe and laugh with joy”. Many of these things, in particular those that have to do with QCD, are of Giorgio’s own making.
Parisi’s contributions are contained in about 40 papers, out of about a hundred he authored during this decade. His other interests included various more formal aspects of quantum field theory, including its first applications to statistical mechanics; the theory and phenomenology of quantum electrodynamics, of great relevance at Frascati, where he worked and where AdA, the first ever collider had been conceived and built, and the ADONE collider started operating in 1969; various aspects of hadron phenomenology and flavor physics [4]; and a constant interest in non-perturbative methods which eventually, towards the end of the decade, developed into an interest in lattice gauge theories that would lead to seminal contributions a few years later [5, 6].
These contributions can be broadly subdivided into three partly overlapping epochs. A first epoch, from 1970 until about 1973, during which the main issue was understanding the meaning of the parton model, explaining its phenomenological success, and understanding its possible theoretical underpinnings and its implications. In other words, understanding Bjorken scaling: the surprising discovery that structure functions, the form factors that describe the structure of the proton as probed by the deep-inelastic scattering of electrons (DIS) are approximately scale dependent. A second epoch, ranging from 1972 to about 1976, in which the nature of QCD as a quantum field theory starts emerging, and the parton model is understood as a limit of perturbative QCD — culminating with the seminal paper [7] that gives its name to this contribution. In other words, understanding scaling violations: the weak logarithmic scale dependence of hadronic structure. And a last epoch, from 1976 to the end of the decade, in which the foundations of perturbative QCD are laid, by introducing some of its basic ideas and developing its fundamental tools.
The vast majority of these contributions are contained in very short papers, with Parisi as a single author: typically two or three pages, in which a seminal idea is presented, in the guise of a flash of light in the dark. The purpose of this brief contribution is to provide an overview of this work, based on the full list of publications of Giorgio Parisi, by discussing most, if not all of the relevant contributions, in an attempt of capturing some of this light.
2 Scaling and the parton model
The earliest Parisi paper on the phenomenology of the strong interactions [8] (and his third published paper on record — the two previous one dealing with quantum field theory topics) concerns what in many current standard textbooks is indeed the first application of QCD to be discussed (see e.g. Refs. [9, 10]): the computation of the total rate of production of hadrons at an electron positron-collider, now usually presented as the ratio
i.e. the ratio of hadron to muon-antimuon production cross-sections. The well-known leading-order result , where are the electric charges of partons, i.e. the constituent of the proton as probed in the perturbative domain, is given as Eq. (1) of this paper.
The derivation is quite close to the modern textbook derivation [9], which is based on the Wilson expansion: indeed, the result is obtained by writing the cross-section in terms of a current matrix element, and evaluating the latter as a vacuum expectation value. The dominance of the contribution from operators with the lowest dimension, that in modern language follows from the Wilson expansion, here instead is justified using an argument that captures the basic underlying physics, namely, the scale hierarchy between the hard annihilation, and the softer dynamics of the hadronization. The argument was previously proposed in the context of DIS [11] and is thereby generalized to annihilation.
The physics question that underlies this paper is made explicit in the title of a companion paper [1]: what is the nature of partons? Are they quarks and other “mythical” (sic) constituents of hadrons? Or are they just hadrons? In the latter case, it is pointed out that DIS would be probing the structure of the vacuum, not of hadrons: the virtual photon would be creating a hadron-antihadron pair out of the vacuum. The paper explores this latter option, and it shows how it could be tested experimentally, by measuring the ratio of longitudinal to transverse cross-sections, which in this case could be determined using SU(3) considerations.
The tension between the parton model viewpoint and the opposite option of computing in terms of hadronic states is bypassed by taking a “duality” approach in Ref. [12], i.e. assuming that results can be equivalently obtained in terms of partons or resonances. Combining this assumption with SU(3) flavor symmetry considerations leads to the sum rule
for the proton-neutron difference of the structure functions that parametrize the deep-inelastic cross-section. Amusingly, in more recent textbooks (see e.g. Ref. [13]) this result is derived using naive parton model argument, and referred to as the Gottfried [14] sum rule — though it is nowhere to be found in Gottfried’s paper [14]. The discovery of its experimental violation in the early nineties arguably opened up the modern era of understanding of the parton substructure of the nucleon [15].
The ultimate underlying question is of course whether Bjorken scaling and the parton model could be justified theoretically. The question is addressed directly in Ref. [16], in which Bjorken scaling is derived using renormalization group arguments, from the Wilson expansion and the Callan-Symanzik equation, under the assumption that the function of the strong interaction Lagrangian has a nontrivial fixed point, where the theory becomes conformal. The derivation is essentially identical to the standard current textbook argument, except that the latter is based on expanding perturbatively in the coupling about zero, so scaling only holds asymptotically and the parton model receives logarithmic corrections — as Parisi’s work would fully elucidate only a few years later.
Before turning to the line of thought that led to this celebrated result, it is worth spending a few words to discuss some studies that were part of investigations of QED, rather than QCD, but are precursors of techniques that are part of the standard QCD toolbox, and indeed some of which played an important role in subsequent QCD studies of Giorgio’s. The first [17] is a discussion of the process in QED in the hard Bremsstrahlung limit, in which the photon energy is large. This paper is based on the use of the Weizsäcker-Williams, or quasi-real photon approximation, in which the photon emission rate is computed in terms of a universal probability for the electron to split into a photon. This approximation will later play an important role in the derivation of evolution equations for parton distributions, namely, the Altarelli-Parisi equations, that I will discuss in Sect. 3 below. In this paper, the result found using this approximation is compared to the exact result. The approximation holds in the limit of collinear photon emission, but it is shown to provide an acceptable approximation for surprisingly large angles — a fact that, with hindsight, explains the unexpected phenomenological success of the Altarelli-Parisi equation, and consequently of Bjorken scaling, even at the relative low scales that were experimentally probed at that time, a phenomenon that was later called precocious scaling.
The second is the companion paper [18], in which the exact computation for this QED process is presented. In this work it is observed that, when computing Feynman diagrams with several particles in the final state, the use of helicity amplitudes instead of traces is highly advantageous. The computation is then performed using helicity amplitude methods, now standard, but only popularized by others many years later [19]. The computational technique itself is in fact the main point of the paper, whose stated goal in its abstract is to “propose a general method to evaluate Feynman graphs” which “avoids the calculation of very long traces”.
3 Scaling violations
Only three months (based on submission dates) elapsed between the proof that Bjorken scaling holds in the presence of a nontrivial fixed point [16], and the first exploration of perturbative scaling violations [20]. Indeed, it had been meanwhile realized that Bjorken scaling is only approximate: deep-inelastic observables are not scale independent, rather, they display a weak logarithmic dependence on the scale at which the target proton is probed, rather that the strong power-like behavior that one might expect based on naive dimensional analysis. The name of the game thus had changed from understanding scaling, to understanding the nature of these logarithmic scaling violations.
The paper [20], with the seemingly modest goal of “showing how to use the data on deep-inelastic scattering to put bounds on the values of anomalous dimensions of the operators involved in the Wilson expansion of the product of two currents near the light-cone”, makes a striking observation. Namely, that if a moment (integral) of a structure functions displays a scaling law
characterized by an anomalous dimension , (where I have used modern notation), then the convolution theorem for Mellin transforms implies that
where the anomalous dimension is the Mellin transform of :
In other words, the finite scale dependence Eq. (3), in which plays the role of an initial condition, emerges as the consequence of an infinitesimal scale transformation from to , given by Eq. (4).
This result is used to get an estimate of the size of the coefficient (which is viewed as a normalization, rather than a coupling), based on an assumed plausible form of the dependence of the anomalous dimension on . The truly astonishing nature of the paper is the fact that Eq. (4) appears as the consequence of the somewhat casual side-remark, that the convolution theorem for Mellin transforms implies that a scaling law of the form of Eq. (3), which is obeyed by moment integrals of the structure function , can be cast in equivalent form as a scaling law satisfied by the structure function itself, governed by a kernel implicitly defined by Eq. (5). At this time, this might appear as a purely mathematical trick, but it effectively contains the basic physics of the Altarelli-Parisi equation, that we will soon get to. Namely, that the kernel — the Altarelli-Parisi splitting function, in modern terminology — provides “the violation in each point of Bjorken scaling law”, i.e. the local scaling violation for each value of , rather than that of the integral of the structure function over .
The question of how to actually compute the anomalous dimension itself is addressed in Ref. [21], in which it is shown that the scaling violations of the moments of structure functions that appear on the left-hand side of Eq. (3) are governed by the anomalous dimensions of the matrix elements of operators with the lowest dimension and increasing spin, i.e. leading-twist operators (again using modern terminology). These are in turn computed in theory: the tour de force calculation is based on solving the Dyson equation satisfied by the operator vertex, rather than using the Wilson expansion as in the modern textbook approach.
A relevant observation is that all these developments predate QCD as a quantum field theory of the strong interaction, and in particular they predate the discovery of the in 1974, the so-called November revolution [22] that convinced the community that QCD is the correct theory of the strong interactions. Their significance became clear once a non-abelian gauge theory was accepted as the quantum field theory of the strong interaction, and it was realized that this theory is amenable to a perturbative treatment in the large-momentum limit.
Indeed, in Ref. [23] the mathematical insight of Ref. [20] and Eq. (4) is combined with the newly acquired knowledge from the seminal work of Gross and Wilczeck [24] on anomalous dimensions and scaling laws of moments of structure functions, to show that structure functions satisfy an integro-differential evolution equation that determines their scale dependence. This equation takes the form (in modern notation)
where is the difference between proton and neutron deep-inelastic structure functions, is an asymptotically vanishing coupling, and is related through Eq. (5) to the nonsinglet anomalous dimension that had been computed by Gross and Wilczek. Amusingly, this paper still predates by several months the discovery of the . Once again, as in Ref. [20], the fact that the scaling law satisfied by moments of structure functions can be equivalently viewed as the consequence of an evolution equation satisfied by the structure function itself, with an evolution kernel related by Mellin transformation to the scaling anomalous dimension, is presented as a mathematical observation. In the words of Ref. [23], the scaling law and the evolution equation “are mathematically equivalent, however we believe that the second equation is easier to test”. In fact, in this work and subsequent investigations based on it [25, 26] the emphasis is on exploiting these results to obtain testable predictions.
However, at this point the step to realizing that this mathematical observation actually has a profound physical meaning — the step leading to the Altarelli-Parisi equation — is very short indeed. A preview is given in the 1976 Moriond lectures, Ref. [27], whose abstract states “the theory of scaling violations in deep inelastic scattering is presented using the parton model language”. In these lectures, evolution equations of the form of Eq. (6) are derived in quantum electrodynamics (QED) for the “constituents of the electron”. Namely, it is observed that an electron can radiate photons that in turn radiate electron-positron pairs, so when probing an electron one is actually probing its internal structure in terms of photons, electrons and positrons. By observing that (at leading perturbative order) the structure function in Eq. (6) is proportional to the number density of electrically charged constituents of the target electron, the evolution equation Eq. (6) is then interpreted as a consequence of the fact that this number density, i.e. the internal structure of the electron, depends on the scale at which the electron is probed because of the scale dependence induced by this radiation process.
The computation is performed using the Weizsäcker-Williams (or equivalent photon) approximation that was already encountered in Ref. [17]. The Weizsäcker-Williams method shows that quantum interference in the radiation process is suppressed by inverse powers of the radiation scale, so, for QED, it is suppressed by powers of , where is the electron mass, and is the scale at which the electron is probed. Consequently the interaction rate — the photon-electron cross-section in this case — factorizes in terms of an elementary photon-electron cross-section, times a universal radiation rate that does not depend on the specific process but rather is a property of the target (the electron, in this case) and can be expressed as a radiation pseudo-probability — not quite a probability because in general it is not positive definite.
It is then observed that if one tries to apply the same reasoning to QCD, namely to the structure of the proton, rather than the electron, there is a fundamental difference, namely, that in QED the vacuum behaves as a dielectric, and consequently the electric charge is screened at large distances, and increases at shorter distances (an argument that is now textbook [9]). In QCD the opposite happens: strongly interacting matter is an “enantion” (loosely translatable as a “contrarian”), it anti-screens, and consequently the charge decreases at shorter distances — asymptotic freedom. Combining these two physical pictures, scale dependence from radiation, and asymptotic freedom from anti-screening, a simple physical picture of the scaling laws of proton structure in QCD emerges. Hence, the key insight in these lectures is to interpret the scaling laws obeyed by deep-inelastic structure function as a consequence of the scaling law of the number densities of partons — the proton’s constituents. This then provides a clear physical picture of scale dependence from evolution equations satisfied by structure functions as a showering process, as realized in modern Monte Carlo parton shower codes [28]. The solution to these evolution equations corresponds to summing a class of Feynman diagrams that had been considered previously by Gribov and Lipatov, who had studied deep-inelastic scattering in a theory with massive vector mesons a few years earlier [29]. However, the derivation presented here is completely independent, and in fact the physical picture of evolution equations is missing in the Gribov-Lipatov approach.
All the pieces of the puzzle come together in the subsequent celebrated, seminal Altarelli-Parisi paper [7] (of which the Moriond lectures [27] are presented as a “preliminary, less complete version”). This is of course one of the most widely cited papers in high-energy physics. In this paper the focus shifts from the deep-inelastic structure function, to the structure of the proton itself. Indeed, the first insight is that the scale dependence of structure function is a reflection of an underlying scale dependence of the structure of the proton, as reflected by parton distribution functions, that provide number densities for the constituents of the proton. The second insight is that the previous mathematical observation [20, 23], that scaling laws can be obtained by solving an integro-differential equations and taking the Mellin transform of the result, can be combined with the physical idea that the proton constituents can be viewed as composite systems, an idea previously suggested by Kogut and Susskind [30].
The combination of these two insights is made quantitative by explicitly computing the probability density for finding a constituent of the proton inside another constituent, i.e. the probability for a parton to split into two. The latter is simply given by the universal rate of emission of an extra final-state parton from an initial state one, which can be computed, as I have already discussed, using the Weizsäcker-Williams approximation, in terms of “splitting functions” that are completely determined by transition matrix elements from the initial parton to the final parton pair. The relevant emission matrix elements are determined using the so-called “old”, i.e. time-ordered perturbation theory, a method that is rarely [31] discussed in modern textbook treatments, and that has been recently revived [32].
The final result is a pair of coupled evolution equations for parton distributions of quarks of the -th flavor or antiflavor and gluons , with , the fraction of the parent proton energy-momentum carried by the given parton, the scale at which the proton is probed, and an arbitrarily chosen reference scale. These have a form that is entirely analogous to that of Eq. (6), namely
where the splitting functions express the probability to find a parton (quark or gluon) inside a parton with fraction of the parent parton’s momentum. These are the celebrated Altarelli-Parisi equations. The same equations are then derived also in the case of polarized parton distributions, which express the probability of the parton to carry the parent proton’s spin.
The derivation of these equations presented in Ref. [7] is arguably rather more transparent than the majority of current available textbook treatments. However, their significance goes well beyond the technical aspects of their derivation. In fact, the fundamental meaning of this work is sometimes not fully appreciated. Indeed, this work is sometimes described [9] as an independent re-derivation of the previous argument by Gribov and Lipatov [29], that shows how to derive evolution equations for deep-inelastic structure functions themselves, rather than their moments. Whereas of course the Altarelli-Parisi paper does contain, as it is stated in its introduction “an alternative derivation of all results of current interest for the behavior of deep inelastic structure functions” (alternative to the derivation based on the scale dependence of matrix elements of leading-twist operators, that is), this is not the main point. Rather, the point, as stated in the title of the paper, is to express quantum-field theoretical results of QCD in parton language. In fact, this work brings to completion the research program and line of thought that started in the early papers that asked the questions: what is the nature of partons? and why does the parton model work?
Indeed, even in modern textbook treatments [10] a parton model language and parton-based approach to QCD factorization is often taken for granted: it is assumed that hard processes can be described by starting with the parton model, and computing perturbative corrections to it. Yet, the theoretical justification for this is unclear — it is unclear why the simple assumptions of the parton model are a reasonable starting point111See e.g. Sect. 9.11 of Ref. [33] for a modern critique of a naive parton-based approach.. What the Altarelli-Parisi paper shows, is that, for deep-inelastic structure functions, a parton-based computation is mathematically equivalent to the result found starting from the Wilson expansion, and then solving renormalization-group equations for Wilson coefficients. Indeed, a terse comment in the conclusion explains “the reasons for the success of this simple method” — the renormalization group equations. So what the Altarelli-Parisi paper amounts to, is a rigorous proof of QCD factorization, albeit in the limited case of deep-inelastic scattering. The paper is at the basis of the modern description of the structure of the proton in terms of parton distributions (see Ref. [34] and ref. therein) but its true significance is in establishing perturbative factorization.
4 The birth of QCD phenomenology
With QCD factorization on a solid field-theoretic basis, at least for deep-inelastic scattering, the leap to full-fledged QCD phenomenology was the logical development. This was laid out in a series of papers, submitted within the short span of two years, which set the basis of QCD calculations for hadronic processes [35, 36], the definition of jet substructure observables [37, 2], the determination of parton distributions [38], the scale dependence of fragmentation functions [39], transverse momentum resummation [40], and threshold resummation [41]. Many of these contributions anticipate ideas of methods that have marked the development of QCD phenomenology, sometimes several decades later.
Some of these papers had a large impact, and shaped subsequent developments, though some were only recognized at a later stage, and some not at all. In particular, Ref. [35] presents a first computation of the dimuon transverse momentum distribution in Drell-Yan production. The celebrated Collins-Soper-Sterman (CSS) work [42], that later set the modern standard for the perturbative calculation for this process, refers to Ref. [35] as the foundation of the whole field: “with the advent of QCD factorization one was able to write the DY cross-section at measured ”. Interestingly, the subsequent paper Ref. [36] considers the effect of taking into account the transverse momentum of partons, a topic that has only become fashionable in this century, with studies of transverse-momentum dependent parton distributions (TMDs).
The leading-order computation of Ref. [35] immediately raised the question of the instability of its result towards the inclusion of logarithmically enhanced perturbative corrections in the small region. The all-order resummation of these terms is accomplished in Ref. [40], by observing that they stem from the emission of soft (i.e. very low energy) gluons, whose contribution was known [43] to exponentiate. Again, in CSS [42] this later paper is described with the words “These authors introduced more powerful techniques: they worked with the Fourier transform with respect to …and they showed the usefulness of soft gluon methods.” The same soft gluon methods, and in particular the powerful idea of performing resummation in terms of a variable which is conjugate by integral transform (Mellin transform, in this case) to the physical variable whose soft limit is being considered, are used in Ref. [41] to perform leading-log threshold resummation of the DIS and DY cross-sections, a result only extended to next-to-leading log about ten years later [44, 45].
In fact, in the brief space of two pages the single-author Ref. [41] manages to contain the following insights, for which essentially no proof is given: the kinematic origin of the different behavior of the DY and DIS cases (proven in detail in Ref. [46]), the impact of running-coupling effects on resummation (fully proven in Ref. [47]), and finally the impact of the analytic continuation from spacelike to timelike in relating DY to DIS, fully studied more than ten years later in Ref. [48], and whose phenomenological consequences were re-discovered in Ref. [49].
Separate lines of developments are related to the needs of precision phenomenology, thereby anticipating the LHC era. In Ref. [38] it is suggested that an accurate parametrization of parton distributions requires extending the standard parametrization that reproduces the limiting behaviors at large and small momentum fractions with bases of orthogonal polynomials, something that groups still using this kind of parametrization first started doing only with the advent of the LHC [50].
In Ref. [37] the energy-momentum correlation of pairs of hadrons in the final state of a collision (hadron-hadron, lepton-hadron, or lepton-lepton) is introduced. This is the precursor of event shape variables [10]: this work is typically cited as the origin of the so-called -parameter, and energy correlators in particular have recently become an extremely fashionable subject (see e.g. Ref. [51] and ref. therein). However, it is interesting to observe that the idea here is somewhat different: namely, to define the observable in the whole (“superinclusive”) final state, rather than inside a jet. This leads to a quantity that can be computed purely using renormalization-group arguments, a suggestion that to the best of our knowledge has never been pursued. A subsequent paper in which a jet structure observable is computed perturbatively [2] is one of only eleven papers out of the 230 published by Parisi to this day to have received no citations.
A similarly limited impact is that of Ref. [39], which presents a first computation of the scale dependence of fragmentation functions. These describe the rate at which partons can fragment into final-state hadrons and are thus the timelike counterpart of the parton distributions which describe the rate at which partons can be pulled out of an initial-state hadron; they are also currently a fashionable topic in view of their future measurement at the forthcoming Electron-Ion Collider (see e.g. Ref. [52] and ref. therein). Their leading-order scale dependence is also governed by the Altarelli-Parisi equation, as this paper first showed. The paper, however, collected only four citations, despite being almost contemporary to the work that is usually presented as the first derivation of this result [53].
5 The work and its context
The last paper by Giorgio Parisi dealing with perturbative QCD is Ref. [41], published nine and a half years after the first [8]. Parisi’s QCD work in its complex is specifically characterized by a striking feature, namely, the desire to understand information coming from experiments, specifically with the language of quantum field theory, and to turn this understanding into testable predictions. Despite their brevity, essentially all of the papers discussed here end with a quantitative estimate, or a brief discussion of phenomenology. Indeed, quantum field theory is the common thread that connects all papers published by Parisi during this decade, ranging from its more formal aspects, to its applications to particle physics and later also statistical mechanics and condensed matter phenomenology, with non-perturbative aspects of field theory providing the other strong connection to phenomenology.
It is indeed the transition from perturbative to non-perturbative, characterized by studies of the divergence of the perturbative expansion in field theory [54], and the nonperturbative study of field theory on the lattice [5, 6], that gradually became the dominant themes of research towards the end of the decade. And indeed in the subsequent decade Parisi’s work shifted mostly to lattice field theory, and also statistical field theory [54, 55, 56], thereby completing the transition from perturbative to non-perturbative of his line of thought.
Within the more general context of research in perturbative QCD, Parisi’s contribution was instrumental in establishing quantum field theory as the paradigm of the theory of fundamental interactions, and quantum chromodynamics as the paradigm of quantum field theory.
Acknowledgments and disclaimer
I am grateful to the organizers of the seminar series “The interdisciplinary contribution of Giorgio Parisi to theoretical physics”, that took place at Sapienza University of Rome in 2022-2023, for inviting me to give a talk in the series, and for subsequently taking the initiative of collecting in a volume the write-ups of all talks. Unlike most of the other contributors, I did not have the privilege to collaborate with Giorgio directly. Somehow, despite my longtime association with my late teacher, mentor and friend Guido Altarelli, I never managed to discuss his joint work with Parisi with him. Moreover, Parisi’s QCD work was done at a time when I was a school kid, so I could not witness its development first-hand. I consequently had to approach it with the attitude of a historian (which I am not), and to some extent of a teacher (which I certainly am). I thus have the obligation of apologizing in advance for any misunderstanding of which I might well have been responsible. Among the organizers, I am especially grateful to Marco Bonvini and Maria Chiara Angelini for comments and criticism on a preliminary version of this manuscript. I would finally like to thank Giorgio for graciously sitting through my talk and sharing with me some comments, thoughts and anecdotes.

### v2

@@
 A first epoch, from 1970 until about 1973, during which the main issue was understanding the meaning of the parton model, explaining its phenomenological success, and understanding its possible theoretical underpinnings and its implications.
-In other words, understanding Bjorken scaling: the surprising discovery that structure functions, the form factors that describe the structure of the proton as probed by the deep-inelastic scattering of electrons (DIS) are approximately scale dependent.
-A second epoch, ranging from 1972 to about 1976, in which the nature of QCD as a quantum field theory starts emerging, and the parton model is understood as a limit of perturbative QCD — culminating with the seminal paper [7] that gives its name to this contribution.
+In other words, understanding Bjorken scaling: the surprising discovery that structure functions, the form factors that describe the structure of the proton as probed by the deep-inelastic scattering of electrons (DIS) are approximately scale independent.
+A second epoch, ranging from 1972 to about 1976, in which the nature of QCD as a quantum field theory starts emerging, and the parton model is understood as a limit of perturbative QCD — culminating with the seminal paper [7] that gives its name to the present paper.
 In other words, understanding scaling violations: the weak logarithmic scale dependence of hadronic structure.
@@
 the ratio of hadron to muon-antimuon production cross-sections.
-The well-known leading-order result , where are the electric charges of partons, i.e.
-the constituent of the proton as probed in the perturbative domain, is given as Eq.
-(1) of this paper.
+Equation (1) of this paper states the now well-known leading-order result: , where are the electric charges of partons, i.e.
+the constituents of the proton as probed in the perturbative domain, and the sum runs over all constituents, and thus can be used to establish both the nature and the total number of constituents — and thus eventually provided evidence for the existence of three quark colors.
 The derivation is quite close to the modern textbook derivation [9], which is based on the Wilson expansion: indeed, the result is obtained by writing the cross-section in terms of a current matrix element, and evaluating the latter as a vacuum expectation value.
@@
 Ref.
-[13]) this result is derived using naive parton model argument, and referred to as the Gottfried [14] sum rule — though it is nowhere to be found in Gottfried’s paper [14].
+[13]) this result is derived using naive parton model arguments, and referred to as the Gottfried [14] sum rule — though it is nowhere to be found in Gottfried’s paper [14].
 The discovery of its experimental violation in the early nineties arguably opened up the modern era of understanding of the parton substructure of the nucleon [15].
@@
 The derivation is essentially identical to the standard current textbook argument, except that the latter is based on expanding perturbatively in the coupling about zero, so scaling only holds asymptotically and the parton model receives logarithmic corrections — as Parisi’s work would fully elucidate only a few years later.
-Before turning to the line of thought that led to this celebrated result, it is worth spending a few words to discuss some studies that were part of investigations of QED, rather than QCD, but are precursors of techniques that are part of the standard QCD toolbox, and indeed some of which played an important role in subsequent QCD studies of Giorgio’s.
+Before turning to the line of thought that led to this celebrated result, it is worth spending a few words to discuss some studies that were part of investigations of QED, rather than QCD, but are precursors of techniques that are now part of the standard QCD toolbox, and indeed some of which played an important role in subsequent QCD studies of Giorgio’s.
 The first [17] is a discussion of the process in QED in the hard Bremsstrahlung limit, in which the photon energy is large.
@@
 In this paper, the result found using this approximation is compared to the exact result.
-The approximation holds in the limit of collinear photon emission, but it is shown to provide an acceptable approximation for surprisingly large angles — a fact that, with hindsight, explains the unexpected phenomenological success of the Altarelli-Parisi equation, and consequently of Bjorken scaling, even at the relative low scales that were experimentally probed at that time, a phenomenon that was later called precocious scaling.
+The approximation holds in the limit of collinear photon emission, but it is shown to provide an acceptable approximation for surprisingly large angles.
+With hindsight, this explains the unexpected phenomenological success of the Altarelli-Parisi equation, and consequently of Bjorken scaling, even at the relative low scales that were experimentally probed at that time, a phenomenon that was later called precocious scaling.
 The second is the companion paper [18], in which the exact computation for this QED process is presented.
-In this work it is observed that, when computing Feynman diagrams with several particles in the final state, the use of helicity amplitudes instead of traces is highly advantageous.
-The computation is then performed using helicity amplitude methods, now standard, but only popularized by others many years later [19].
+In this work it is observed that, when computing Feynman diagrams with several particles in the final state, it is highly advantageous to use, instead of Dirac traces, the helicity amplitude method, previously introduced in Ref.
+[19].
+The computation is then performed using this method, now standard, but only popularized by others many years later [20].
 The computational technique itself is in fact the main point of the paper, whose stated goal in its abstract is to “propose a general method to evaluate Feynman graphs” which “avoids the calculation of very long traces”.
@@
 It is then observed that if one tries to apply the same reasoning to QCD, namely to the structure of the proton, rather than the electron, there is a fundamental difference, namely, that in QED the vacuum behaves as a dielectric, and consequently the electric charge is screened at large distances, and increases at shorter distances (an argument that is now textbook [9]).
-In QCD the opposite happens: strongly interacting matter is an “enantion” (loosely translatable as a “contrarian”), it anti-screens, and consequently the charge decreases at shorter distances — asymptotic freedom.
+In QCD the opposite happens: strongly interacting matter is, in the words of Ref.
+[28], an “enantion”, loosely translatable as a “contrarian”: it anti-screens, and consequently the charge decreases at shorter distances — asymptotic freedom.
 Combining these two physical pictures, scale dependence from radiation, and asymptotic freedom from anti-screening, a simple physical picture of the scaling laws of proton structure in QCD emerges.
@@
 In this paper the focus shifts from the deep-inelastic structure function, to the structure of the proton itself.
-Indeed, the first insight is that the scale dependence of structure function is a reflection of an underlying scale dependence of the structure of the proton, as reflected by parton distribution functions, that provide number densities for the constituents of the proton.
+Indeed, the first insight is that the scale dependence of structure function is a consequence of an underlying scale dependence of the structure of the proton, as reflected by parton distribution functions, that provide number densities for the constituents of the proton.
 The second insight is that the previous mathematical observation [21, 24], that scaling laws can be obtained by solving an integro-differential equations and taking the Mellin transform of the result, can be combined with the physical idea that the proton constituents can be viewed as composite systems, an idea previously suggested by Kogut and Susskind [31].
@@
 However, their significance goes well beyond the technical aspects of their derivation.
-In fact, the fundamental meaning of this work is sometimes not fully appreciated.
-Indeed, this work is sometimes described [9] as an independent re-derivation of the previous argument by Gribov and Lipatov [29], that shows how to derive evolution equations for deep-inelastic structure functions themselves, rather than their moments.
+In fact, the fundamental meaning of this paper is sometimes not fully appreciated.
+Indeed, this work is sometimes described [9] as an independent re-derivation of the previous argument by Gribov and Lipatov [30], that shows how to obtain evolution equations for deep-inelastic structure functions themselves, rather than their moments.
 Whereas of course the Altarelli-Parisi paper does contain, as it is stated in its introduction “an alternative derivation of all results of current interest for the behavior of deep inelastic structure functions” (alternative to the derivation based on the scale dependence of matrix elements of leading-twist operators, that is), this is not the main point.
@@
 So what the Altarelli-Parisi paper amounts to, is a rigorous proof of QCD factorization, albeit in the limited case of deep-inelastic scattering.
-The paper is at the basis of the modern description of the structure of the proton in terms of parton distributions (see Ref.
+The paper is at the basis of the modern description of the structure of the proton in terms of parton distributions (see e.g.
+Ref.
 [35] and ref.
@@
 This was laid out in a series of papers, submitted within the short span of two years, which set the basis of QCD calculations for hadronic processes [36, 37], the definition of jet substructure observables [38, 2], the determination of parton distributions [39], the scale dependence of fragmentation functions [40], transverse momentum resummation [41], and threshold resummation [42].
-Many of these contributions anticipate ideas of methods that have marked the development of QCD phenomenology, sometimes several decades later.
+Many of these contributions anticipate ideas and methods that have marked the development of QCD phenomenology, sometimes several decades later.
 Some of these papers had a large impact, and shaped subsequent developments, though some were only recognized at a later stage, and some not at all.
@@
 Interestingly, the subsequent paper Ref.
-[36] considers the effect of taking into account the transverse momentum of partons, a topic that has only become fashionable in this century, with studies of transverse-momentum dependent parton distributions (TMDs).
+[37] considers the effect of taking into account the transverse momentum of partons, a topic that has only become fashionable in this century, with studies of transverse-momentum dependent parton distributions (TMDs [44]).
 The leading-order computation of Ref.
@@
 I am grateful to the organizers of the seminar series “The interdisciplinary contribution of Giorgio Parisi to theoretical physics”, that took place at Sapienza University of Rome in 2022-2023, for inviting me to give a talk in the series, and for subsequently taking the initiative of collecting in a volume the write-ups of all talks.
-Unlike most of the other contributors, I did not have the privilege to collaborate with Giorgio directly.
+Unlike most of the other contributors, I did not have the privilege of collaborating with Giorgio directly.
 Somehow, despite my longtime association with my late teacher, mentor and friend Guido Altarelli, I never managed to discuss his joint work with Parisi with him.
@@
 Among the organizers, I am especially grateful to Marco Bonvini and Maria Chiara Angelini for comments and criticism on a preliminary version of this manuscript.
+I also thank Giovanni Stagnitto for spotting several typos.
 I would finally like to thank Giorgio for graciously sitting through my talk and sharing with me some comments, thoughts and anecdotes.